OpenStudy (fionacndg):

The length of a rectangle is 20 in and the width is 35 in. The rectangle is then dilated by a scale factor of one-third. What are the new dimensions?

jimthompson5910 (jim_thompson5910):

"dilated by a scale factor of one-third" means each dimension is divided by 3

OpenStudy (bradely):

scale factor (1/3)means multiply with 1/3 length =20*(1/3) width=35*(1/3) Source: http://www.mathskey.com/question2answer/